# Mirror Online Conformal Prediction with Intermittent Feedback

### v1

@@ II-C Bayesian Adaptive Conformal Inference
In Appendix A, we show that IB-ACI achieves sub-linear expected regret (see (14)).
@@ III Intermittent Mirror Online Conformal Prediction
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$. In Appendix B, we show that the regularizing function (9) is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\sigma$.
@@ III-A Theoretical Guarantees
We now analyze the performance of IM-OCP, showing that it satisfies expected long-term coverage while also enjoying sublinear expected regret. Proofs are deferred to Appendix C.
@@ Appendix A Regret Bound of IB-ACI
$r_{t}=\arg\min_{r}h_{t}\psi(r)+\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}.$ (A-1)
@@
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (A-2)
@@
$\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\ell_{1-\alpha}(r_{i+1},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (A-3)
@@
By combining (A-2) and (A-3), we obtain
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (A-4)
@@
$\mathrm{Reg}(T)$ $\leq\mathbb{E}_{\text{obs}_{i}\sim p_{i}}\left[h_{t}\psi(u)+\sum_{i=1}^{T}\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}\right]$ (A-5)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{i}h_{i}\epsilon}$
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}h_{i}\epsilon},$
@@
$\mathrm{Reg}(T)$ $\leq\mathcal{O}\left(\psi(u)\sqrt{T}+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}\epsilon\sqrt{t}})\right)\sim\mathcal{O}(\sqrt{T}).$ (A-6)
Appendix B Proof of Strongly Convex and Smoothness
@@
$R(r)=$ $\mathbb{E}_{r_{i}^{*}\sim P}[\ell_{1-\alpha}(r,r_{i}^{*})]+\frac{\sigma}{2}|r|^{2}$ (B-1a)
$=$ $\int_{0}^{B}\ell_{1-\alpha}(r,r^{*})P(r^{*})\text{d}r^{*}+\frac{\sigma}{2}|r|^{2}$ (B-1b)
$=$ $(\alpha-1)\int_{r}^{B}(r-r^{*})P(r^{*})\text{d}r^{*}$
$+\alpha\int_{0}^{r}(r-r^{*})P(r^{*})\text{d}r^{*}+\frac{\sigma}{2}|r|^{2}.$ (B-1c)
@@
$\nabla R(r)=$ $\alpha\int_{0}^{r}P(r^{*})\text{d}r^{*}+(\alpha-1)\int_{r}^{B}P(r^{*})\text{d}r^{*}+\sigma r$ (B-2)
$=$ $\alpha-\int_{r}^{B}P(r^{*})\text{d}r^{*}+\sigma r$
$=$ $\int_{0}^{r}P(r^{*})\text{d}r^{*}-(1-\alpha)+\sigma r$
$\mathop{=}\limits^{(a)}$$\left\{\begin{array}[]{ll}\mathrm{CDF}_{P}(r)-(1-\alpha)+\sigma r,&0\leq r\leq B\\ -(1-\alpha)+\sigma r,&r<0\\ \alpha+\sigma r,&r>B\end{array}\right.$
@@
$\nabla^{2}R(r)=\left\{\begin{array}[]{ll}P(r)+\sigma,&0\leq r\leq B\\ \sigma,&\text{otherwise}\end{array}\right.$ (B-3)
@@
Appendix C Proof of Theoretical Guarantees in Sec. III-A
In this appendix, we provide a detailed derivation of the theoretical guarantees presented in Sec. III-A. We begin by deriving the bound on the iterates, then establish the miscoverage guarantee and regret bound, and finally discuss the impact of the learning rate on the miscoverage guarantee and regret.
C-A Proof of Lemma 1
@@
$\nabla R(r_{t})-\nabla R(r_{t-1})=-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}.$ (C-1)
Having assumed the mirror map to be $\mu$-strongly convex, it holds
$\left\langle\nabla R(r_{t})-\nabla R(r_{t-1}),r_{t}-r_{t-1}\right\rangle$ (C-2)
$\qquad\qquad\qquad\qquad\qquad\geq\mu\left\|r_{t}-r_{t-1}\right\|^{2}$
and
$\left\langle\nabla R(r_{t})-\nabla R(r_{t-1}),r_{t}-r_{t-1}\right\rangle$ (C-3)
$\qquad\qquad\qquad\leq\frac{1}{\mu}\left\|\nabla R(r_{t})-\nabla R(r_{t-1})\right\|^{2}.$
@@
$r_{t}\leq r_{t-1}-\frac{\eta_{t-1}}{\mu}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}},$ (C-4)
from which we conclude that $E_{t-1}=1$ and $\text{obs}_{t-1}=1$ in order for the inequality $r_{t}>r_{t-1}$ be true. At the same time, given that $E_{t-1}=1$ and $\text{obs}_{t-1}=1$, inequality (LABEL:eq:sec_ineq_sc) implies that
$r_{t-1}\geq r_{t}+\frac{\eta_{t-1}(\alpha-1)}{\mu p_{t-1}}>B,$ (C-5)
@@ C-B Proof of Theorem 1
$\mathbb{E}[\nabla R(r_{t+1})]$ (C-6)
$=\mathbb{E}[\nabla R(r_{\tau})]-\mathbb{E}\left[\sum^{T}_{t=\tau}\eta_{t}(E_{t}-\alpha)\frac{\text{obs}_{t}}{p_{t}}\right]$
$=\mathbb{E}[\nabla R(r_{\tau})]-\sum^{T}_{t=\tau}\eta_{t}(\mathbb{E}[E_{t}]-\alpha).$
@@
$\overline{\mathrm{MisCov}}(T)$ (C-7)
$=\left|\mathbb{E}\left[\frac{1}{T}\sum_{i=1}^{T}\left(\sum_{\tau=1}^{i}\Delta_{\tau}\right)\eta_{i}(E_{i}-\alpha)\right]\right|$
@@
$\leq\frac{1}{T}\sum_{\tau=1}^{T}|\Delta_{\tau}|\cdot\left|\mathbb{E}\left[\sum_{i=\tau}^{T}\eta_{i}(E_{i}-\alpha)\right]\right|$
$\leq\frac{\left\|\Delta_{1:T}\right\|_{1}}{T}\cdot\max_{\tau\in\{1,\dots,T\}}\left|\sum_{i=\tau}^{T}\eta_{i}(\mathbb{E}\left[E_{i}\right]-\alpha)\right|.$
From (C-6), it holds that
$\sum_{i=\tau}^{T}\eta_{i}(\mathbb{E}\left[E_{i}\right]-\alpha)=\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right],$ (C-8)
which allows us to rewrite (C-7) as
$\overline{\mathrm{MisCov}}(T)$ (C-9)
$\leq\frac{\left\|\Delta_{1:T}\right\|_{1}}{T}\max_{\tau\in\{1,\dots,T\}}\left|\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right]\right|.$
@@
$\max_{\tau\in\{1,\dots,T\}}$ $\left|\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right]\right|$ (C-10)
$\leq\max_{\tau\in\{1,2,\dots,T\}}L\left|\mathbb{E}\left[r_{T+1}-r_{\tau}\right]\right|$
$\leq LB+\frac{L\eta_{1}}{\mu p_{\min}},$
@@
By plugging (C-10) into (C-9) and noting that $\left\|\Delta_{1:T}\right\|_{1}=1/\eta_{T}$, we achieve the final result
$\overline{\mathrm{MisCov}}(T)\leq\frac{1}{T\eta_{T}}\left(LB+\frac{L\eta_{1}}{\mu p_{\min}}\right).$ (C-11)
C-C Proof of Theorem 2
@@
$\sum_{t=1}^{T}{\ell_{1-\alpha}(r_{t},r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}}-\sum_{t=1}^{T}{\ell_{1-\alpha}(u,r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}}$ (C-12)
@@
$\mathbb{E}\left[\left|\nabla\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}\right|\right]\leq\mathbb{E}\left[\left|\frac{\text{obs}_{i}}{p_{i}}\right|\right]=1,$ (C-13)
@@
$\overline{\mathrm{Reg}}(T)$ (C-14)
$=\mathbb{E}\left[\sum_{i=1}^{T}{\ell_{\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}}\right]-\sum_{i=1}^{T}{\ell_{\alpha}(u,r_{i}^{*})}$
@@ C-D Proof of Corollary 1
$\overline{\mathrm{MisCov}}(T)\leq$ $\frac{LB}{c}T^{\beta-1}+\frac{L}{c\mu p_{\min}}T^{-1}$ (C-15)
$\sim\mathcal{O}(AT^{\beta-1}),$
@@
$\overline{\mathrm{Reg}}(T)$ $\leq\frac{D_{T}}{c}T^{\beta}+\frac{c}{2\mu p_{\min}}T^{1-\beta}$ (C-16)
$\sim\mathcal{O}(T^{\beta}+T^{\beta-1})=\mathcal{O}(T^{\max\{\beta,\beta-1\}}).$
@@
$\overline{\mathrm{MisCov}}(T)\leq\frac{LB\mu{p_{\min}}+Lc}{\mu p_{\min}c}{T^{\beta-1}}\sim\mathcal{O}(AT^{\beta-1}).$ (C-17)
@@
$\overline{\mathrm{Reg}}(T)$ $=\frac{D_{T}T^{\beta}}{c}+\frac{c}{2\mu p_{\min}}\sum_{t=1}^{T}t^{-\beta}$ (C-18)
$\sim\mathcal{O}(T^{\beta}+T^{1-\beta})=\mathcal{O}(T^{\max\{\beta,\beta-1\}}).$

### v2

@@ II-C Bayesian Adaptive Conformal Inference
In Appendix -A, we show that IB-ACI achieves sub-linear expected regret (see (14)).
@@ III Intermittent Mirror Online Conformal Prediction
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$. In Appendix -B, we show that the regularizing function (9) is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\sigma$.
@@ III-A Theoretical Guarantees
We now analyze the performance of IM-OCP, showing that it satisfies expected long-term coverage while also enjoying sublinear expected regret. Proofs are deferred to Supplementary Material.
@@ V Conclusion
$r_{t}=\arg\min_{r}h_{t}\psi(r)+\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}.$ (16)
@@
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (17)
@@
$\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\ell_{1-\alpha}(r_{i+1},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (18)
@@
By combining (17) and (18), we obtain
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (19)
@@
$\mathrm{Reg}(T)\leq\mathbb{E}_{\text{obs}_{i}\sim p_{i}}\left[h_{t}\psi(u)+\sum_{i=1}^{T}\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}\right]$ (20)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{i}h_{i}\epsilon}\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}h_{i}\epsilon},$
@@
$\mathrm{Reg}(T)$ $\leq\mathcal{O}\left(\psi(u)\sqrt{T}+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}\epsilon\sqrt{t}})\right)\sim\mathcal{O}(\sqrt{T}).$ (21)
-B Proof of Strongly Convex and Smoothness
@@
$R(r)=\mathbb{E}_{r_{i}^{*}\sim P}[\ell_{1-\alpha}(r,r_{i}^{*})]+\frac{\sigma}{2}|r|^{2}$
$=\int_{0}^{B}\ell_{1-\alpha}(r,r^{*})P(r^{*})\text{d}r^{*}+\frac{\sigma}{2}|r|^{2}$ (22)
$=\alpha\int_{0}^{B}(r-r^{*})P(r^{*})\text{d}r^{*}-\int_{r}^{B}(r-r^{*})P(r^{*})\text{d}r^{*}+\frac{\sigma}{2}|r|^{2}.$
@@
$\nabla R(r)=\int_{0}^{r}P(r^{*})\text{d}r^{*}-(1-\alpha)+\sigma r$ (23)
$\mathop{=}\limits^{(a)}\left\{\begin{array}[]{ll}\mathrm{CDF}_{P}(r)-(1-\alpha)+\sigma r,&0\leq r\leq B,\\ -(1-\alpha)+\sigma r,&r<0,\\ \alpha+\sigma r,&r>B.\end{array}\right.$
@@
$\nabla^{2}R(r)=\left\{\begin{array}[]{ll}P(r)+\sigma,&0\leq r\leq B,\\ \sigma,&\text{otherwise}.\end{array}\right.$ (24)
@@
Supplementary Material (Appendix)
-C Proof of Lemma 1
@@
$\nabla R(r_{t})-\nabla R(r_{t-1})=-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}.$ (A-1)
Having assumed the $R(r)$ to be $\mu$-strongly convex, it holds
$\left\langle\nabla R(r_{t})-\nabla R(r_{t-1}),r_{t}-r_{t-1}\right\rangle\geq\mu\left\|r_{t}-r_{t-1}\right\|^{2},$ (A-2)
$\left\langle\nabla R(r_{t})-\nabla R(r_{t-1}),r_{t}-r_{t-1}\right\rangle\leq\frac{\left\|\nabla R(r_{t})-\nabla R(r_{t-1})\right\|^{2}}{\mu}$ (A-3)
@@
$r_{t}\leq r_{t-1}-\frac{\eta_{t-1}}{\mu}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}},$ (A-4)
from which we conclude that $E_{t-1}=1$ and $\text{obs}_{t-1}=1$ in order for the inequality $r_{t}>r_{t-1}$ be true. At the same time, given that $E_{t-1}=1$ and $\text{obs}_{t-1}=1$, inequality (LABEL:eq:sec_ineq_sc) implies
$r_{t-1}\geq r_{t}+\frac{\eta_{t-1}(\alpha-1)}{\mu p_{t-1}}>B,$ (A-5)
@@
$\mathbb{E}[\nabla R(r_{t+1})]$ $=\mathbb{E}[\nabla R(r_{\tau})]-\mathbb{E}\left[\sum^{T}_{t=\tau}\eta_{t}(E_{t}-\alpha)\frac{\text{obs}_{t}}{p_{t}}\right]$
$=\mathbb{E}[\nabla R(r_{\tau})]-\sum^{T}_{t=\tau}\eta_{t}(\mathbb{E}[E_{t}]-\alpha).$ (B-1)
@@
$\overline{\mathrm{MisCov}}(T)$ $=\left|\mathbb{E}\left[\frac{1}{T}\sum_{i=1}^{T}\left(\sum_{\tau=1}^{i}\Delta_{\tau}\right)\eta_{i}(E_{i}-\alpha)\right]\right|$ (B-2)
@@
$\leq\frac{\sum_{\tau=1}^{T}|\Delta_{\tau}|}{T}\cdot\left|\mathbb{E}\left[\sum_{i=\tau}^{T}\eta_{i}(E_{i}-\alpha)\right]\right|$
$\leq\frac{\left\|\Delta_{1:T}\right\|_{1}}{T}\cdot\max_{\tau\in\{1,\dots,T\}}\left|\sum_{i=\tau}^{T}\eta_{i}(\mathbb{E}\left[E_{i}\right]-\alpha)\right|$
From (B-1), it holds that
$\sum_{i=\tau}^{T}\eta_{i}(\mathbb{E}\left[E_{i}\right]-\alpha)=\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right],$ (B-3)
which allows us to rewrite (B-2) as
$\overline{\mathrm{MisCov}}(T)\leq\frac{\left\|\Delta_{1:T}\right\|_{1}}{T}\max_{\tau\in\{1,\dots,T\}}\left|\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right]\right|.$ (B-4)
@@
$\max_{\tau\in\{1,\dots,T\}}\left|\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right]\right|$ (B-5)
$\quad\leq\max_{\tau\in\{1,2,\dots,T\}}L\left|\mathbb{E}\left[r_{T+1}-r_{\tau}\right]\right|LB+\frac{L\eta_{1}}{\mu p_{\min}},$
@@
By plugging (B-5) into (B-4) and noting that $\left\|\Delta_{1:T}\right\|_{1}=1/\eta_{T}$, we achieve the final result
$\overline{\mathrm{MisCov}}(T)\leq\frac{1}{T\eta_{T}}\left(LB+\frac{L\eta_{1}}{\mu p_{\min}}\right).$ (B-6)
-E Proof of Theorem 2
@@
$\sum_{t=1}^{T}{\ell_{1-\alpha}(r_{t},r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}}-\sum_{t=1}^{T}{\ell_{1-\alpha}(u,r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}}$ (C-1)
@@
$\mathbb{E}\left[\left|\nabla\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}\right|\right]\leq\mathbb{E}\left[\left|\frac{\text{obs}_{i}}{p_{i}}\right|\right]=1,$ (C-2)
@@
$\overline{\mathrm{Reg}}(T)=\mathbb{E}\left[\sum_{i=1}^{T}{\ell_{\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}}\right]-\sum_{i=1}^{T}{\ell_{\alpha}(u,r_{i}^{*})}$ (C-3)
@@
$\overline{\mathrm{MisCov}}(T)\leq$ $\frac{LB}{c}T^{\beta-1}+\frac{L}{c\mu p_{\min}}T^{-1}\sim\mathcal{O}(AT^{\beta-1}),$ (D-1)
@@
$\overline{\mathrm{Reg}}(T)$ $\leq\frac{D_{T}}{c}T^{\beta}+\frac{c}{2\mu p_{\min}}T^{1-\beta}\sim=\mathcal{O}(T^{\max\{\beta,\beta-1\}}).$ (D-2)
@@
$\overline{\mathrm{MisCov}}(T)\leq\frac{LB\mu{p_{\min}}+Lc}{\mu p_{\min}c}{T^{\beta-1}}\sim\mathcal{O}(AT^{\beta-1}).$ (D-3)
@@
$\overline{\mathrm{Reg}}(T)$ $=\frac{D_{T}T^{\beta}}{c}+\frac{c}{2\mu p_{\min}}\sum_{t=1}^{T}t^{-\beta}\sim\mathcal{O}(T^{\max\{\beta,\beta-1\}}).$ (D-4)